# Inference-Time Policy Steering through Human Interactions

### v1

@@ II-A Steering Towards User Intent
which captures the average $\mathcal{L}_{2}$-distance between all states in the generated trajectory and the target 3D state $\mathbf{z}$. This objective function allows users to flexibly point goals in a scene, by specifying which objects to manipulate in a real-world kitchen environment (Figure 1).
@@
When the sketch $\mathbf{z}$ has a different length than generated trajectories $\tau$, we uniformly resampled $\mathbf{z}^{\text{sketch}}$ to match the temporal dimension of generated samples. In comparison to the point input, this objective function allows users to specify shape preferences of a trajectory through a directional path in a robot’s workspace (Figure 6).
@@
Figure 4: Alignment vs. Collision in Maze2D. We compare various sampling methods with ACT and DP steered using sketch input. (1) Steering improves alignment at the cost of constraint satisfaction and increased collisions. Moreover, (2) Multimodal policies (DP) steered with PR enhance alignment without significant distribution shift, while (3) unimodal policies (ACT) are harder to steer effectively, particularly if they lack robustness (see Figure 6). Finally, DP steered with SS achieves the best alignment-constraint satisfaction trade-off.
@@ II-B Inference-Time Interaction-Conditioned Sampling
Biased Initialization (BI). In Biased Initialization, inspired by [23], we modify the initialization of the reverse diffusion process. Instead of initializing with a noise trajectory $\tau_{N}$ [^2] Subscript denotes diffusion steps for $\tau_{i}$ and trajectory timesteps for $s_{t}$. $\sim\mathcal{N}(0,I)$, we use a Gaussian-corrupted version of the user input $\mathbf{z^{\text{point}}}$ or $\mathbf{z^{\text{sketch}}}$ as $\tau_{N}$, bringing the process closer to the desired mode from the outset [^3] BI requires DDIM [24] rather than DDPM [25] for reverse diffusion.. While this approach specifies user intent at initialization, the sampling process may still deviate from this input.
@@
where $\epsilon_{\theta}(\tau_{i},i)$ is the denoising network, $\eta\sim\mathcal{N}(0,I)$ is Gaussian noise, $\beta_{i}$ is the guide ratio that controls the alignment gradient’s influence, $\alpha_{i}$, $\gamma_{i}$, $\sigma_{i}$ are diffusion-specific hyperparameters. This alignment gradient steers the reverse process toward trajectories aligned with $\mathbf{z}$, potentially discovering new behavior modes in states where unconditional predictions would otherwise be unimodal and far from $\mathbf{z}$. However, sampling with a weighted sum of denoising and alignment gradients in Equation 5 approximates sampling from the unnormalized weighted sum of the policy distribution and the objective function rather than their product [21], which can result in out-of-distribution samples (Figure 3).
@@
Stochastic Sampling (SS). Finally, in Stochastic Samplingwe use annealed MCMC to optimize the composition of the diffusion model $\pi_{\theta}$ and the objective $\xi(\tau_{i},\mathbf{z})$ [21]. Here, the denoising function $\epsilon_{\theta}(\tau_{i},i)$ at each timestep $i$ represents the score $\nabla_{\tau}\log p_{i}(\tau)$ for a sequence of probability distributions $\{p_{i}(\tau)\}_{0\leq i\leq N}$, where $p_{N}(\tau)$ is Gaussian and $p_{0}(\tau)$ is the distribution of valid trajectories in the environment. Simultaneously, the objective $\xi(\tau,\mathbf{z})$ defines an energy-based model (EBM) distribution $q(\tau)\propto e^{-\xi(\tau,\mathbf{z})}$. Steering toward user intent then corresponds to sequentially sampling from $p_{N}(\tau)q(\tau)$ to $p_{0}(\tau)q(\tau)$, yielding final samples from $p_{0}(\tau)q(\tau)$ that are both valid within the environment and minimize the specified objective.
@@ V CONCLUSION
In this work, we propose the Inference-Time Policy Steering (ITPS) framework, which integrates real-time human interactions to control policy behaviors during inference without requiring explicit policy training. We demonstrate how this approach enables humans to steer policies and benchmark several algorithms across both simulation and real-world experiments. One limitation of our work is the reliance on an expensive sampling procedure to produce behaviors aligned with human intent. In future work, we aim to distill the steering process into an interaction-conditioned policy to achieve faster responses to human interactions and conduct a user study to further validate steerability. We hope this work sheds light on the integration of human interaction with learned generative policies.

### v2

@@ II-A Steering Towards User Intent
which captures the average $\mathcal{L}_{2}$-distance between all states in the generated trajectory and the target 3D state $\mathbf{z}$ [^2] While $\min\limits_{s_{1}\dots s_{T}}\|\mathbf{s}_{t}-\mathbf{z}\|_{2}$ is more accurate, gradients are not smooth.. This objective function allows users to flexibly point goals in a scene, by specifying which objects to manipulate in a real-world kitchen environment (Figure 1).
@@
[…]
@@
Figure 4: Alignment vs. Collision in Maze2D. We compare various sampling methods with ACT and DP steered using sketch input. (1) Steering frozen policies improves alignment at the cost of constraint satisfaction and increased collisions. Moreover, (2) Multimodal policies (DP) steered with PR enhance alignment without significant distribution shift, while (3) unimodal policies (ACT) are harder to steer effectively, particularly if they lack robustness (see Figure 6). (4) Finally, DP steered with SS achieves the best alignment-constraint satisfaction trade-off.
@@ II-B Inference-Time Interaction-Conditioned Sampling
Biased Initialization (BI). In Biased Initialization, inspired by [24], we modify the initialization of the reverse diffusion process. Instead of initializing with a noise trajectory $\tau_{N}$ [^4] Subscript denotes diffusion steps for $\tau_{i}$ and trajectory timesteps for $s_{t}$. $\sim\mathcal{N}(0,I)$, we use a Gaussian-corrupted version of the user input $\mathbf{z^{\text{point}}}$ or $\mathbf{z^{\text{sketch}}}$ as $\tau_{N}$, bringing the process closer to the desired mode from the outset. While this approach specifies user intent at initialization, the sampling process may still deviate from this input.
@@
where $\epsilon_{\theta}(\tau_{i},i)$ is the denoising network, $\eta\sim\mathcal{N}(0,I)$ is Gaussian noise, $\beta_{i}$ is the guide ratio that controls the alignment gradient’s influence, $\alpha_{i}$, $\gamma_{i}$, $\sigma_{i}$ are diffusion-specific hyperparameters. This alignment gradient steers the reverse process toward trajectories aligned with $\mathbf{z}$, potentially discovering new behavior modes in states where unconditional predictions would otherwise be unimodal and far from $\mathbf{z}$. However, sampling with a weighted sum of denoising and alignment gradients in Equation 5 approximates sampling from the weighted sum of the policy distribution and the objective distribution rather than their product [21], which can result in out-of-distribution samples (Figure 3).
@@
Stochastic Sampling (SS). Finally, in Stochastic Sampling, we use annealed MCMC to optimize the composition of the diffusion model $\pi_{\theta}$ and the objective $\xi(\tau_{i},\mathbf{z})$ [21]. Here, the denoising function $\epsilon_{\theta}(\tau_{i},i)$ at each timestep $i$ represents the score $\nabla_{\tau}\log p_{i}(\tau)$ for a sequence of probability distributions $\{p_{i}(\tau)\}_{0\leq i\leq N}$, where $p_{N}(\tau)$ is Gaussian and $p_{0}(\tau)$ is the distribution of valid trajectories in the environment. Simultaneously, the objective $\xi(\tau,\mathbf{z})$ defines an energy-based model (EBM) distribution $q(\tau)\propto e^{-\xi(\tau,\mathbf{z})}$. Steering toward user intent then corresponds to sequentially sampling from $p_{N}(\tau)q(\tau)$ to $p_{0}(\tau)q(\tau)$, yielding final samples from $p_{0}(\tau)q(\tau)$ that are both valid within the environment and minimize the specified objective.
@@ V CONCLUSION
In this work, we propose the Inference-Time Policy Steering (ITPS) framework, which integrates real-time human interactions to control policy behaviors during inference without requiring explicit policy training. We demonstrate how this approach enables humans to steer policies and benchmark several algorithms across both simulation and real-world experiments. One limitation of our work is the reliance on an expensive sampling procedure to produce behaviors aligned with human intent. In future work, we aim to distill the steering process into an interaction-conditioned policy to achieve faster responses to human interactions and conduct a user study to further validate steerability.